NvidiaIntely (nvidiaintely):

A common stain remover is sold in concentrated form as 1.8M sodium hypochlorite (NaOCl). For everyday use, you create 2.2 liter of a more dilute solution by dissolving 0.015 liters of the 1.8M solution in water. What is the concentration of this new solution? 0.012M 0.18M 1.4M 2.64M

NvidiaIntely (nvidiaintely):

@sweetburger @welshfella

NvidiaIntely (nvidiaintely):

@Nnesha @TheSmartOne

OpenStudy (photon336):

\[n_{a}c_{a} = n_{b}c_{b}\] where n = concentration c = volume

OpenStudy (photon336):

\[n_{a} = 1.8M; 0.15L = c_{a}; 2.2L = c_{b}\] \[\frac{ n_{a}c_{a} }{ c_{b} } = n_{b}\] \[\frac{ 1.8~\frac{ mol }{ L }*0.15~L }{ 2.2~L } = c_{b}\]